If 6.5 gram of magnesium ribbon reacts completely with hydrochloric acid how many grams of H2 will be liberate (H=1,Mg=24)


1
Expert's answer
2022-02-10T19:58:02-0500

The reaction that occurs is

"Mg_{(s)}+2HCl_{(ac)} \to H_{2(g)}+MgCl_{2(ac)}"

We use the mass of Mg to find how much H2 was produced

"\text{6.5 g Mg}\,(\frac{\text{1 mol Mg}}{\text{24.0 g Mg}})(\frac{\text{1 mol H}_2}{\text{1 mol Mg}})(\frac{\text{2 g H}_2}{\text{1 mol H}_2})
\\ = \text{0.5417 g H}_2"


In conclusion, 0.5417 g of H2(g) are produced.